OpenStudy (anonymous):

A (0,0) B(6,3) ½ of the way from A to B

OpenStudy (anonymous):

vectors

OpenStudy (jit4won):

you can solve it without using vectors

OpenStudy (jit4won):

if c be the center of AB coordinates of C = ( 0+6/2 , 0+3 /2) = (3,1.5) using distance finding formula we get AC=\[\sqrt{(3-0)^2 + (1.5-0 ) ^2}\]

OpenStudy (jit4won):

3.354101966

OpenStudy (jit4won):

in vectors you may write 3 i +1.5 j